OpenStudy (anonymous):

find the slope of the line through the pair of points. (–1, –6) and (1, 12)

OpenStudy (anonymous):

Rise over run. Rise is: \[y_2 - y_1\] Run is: \[x_2 - x_1\] So the slope is: \[\frac{y_2-y_1}{x_2-x_1}\] We know: \[x_1 = -1,\ y_1 = 6,\ x_2 =1,\ y_2=12\ \] Can you plug them into the formula?

OpenStudy (anonymous):

m=18/2 or m=9